OpenStudy (anonymous):

The area of a rectangle is measured by the function A(x) = 60x − x2, where x is the length of the rectangle and is measured in yards. (a) Evaluate A(2) and explain what the dependent and independent variables represent. (b) What is the area of the rectangle when its length is 6 yards? (c) What is the rectangle’s length when its area is 800 square yards?

OpenStudy (anonymous):

\[A(x)=60x-x^2\] A(4)=(60(4)-16)=240-16=242 The dependent Variable is Area and The independent is the length,a. b)A(6)=60(6)-36=324yards c)800=60x-x^2=-x^2+60x-800=x=20